Please help me solve this problem.



A coin is tossed 3 times. What is the probability of obtaining at least 1 head?

Found 2 solutions by stanbon, subudear:
Answer by stanbon(75887)

P(x>= 1) = 1-P(no heads in three tosses)
= 1- (1/2)^3
= 1 - (1/8)
= 7/8
===========
Cheers,
Stan H.

Answer by subudear(62)

A coin is tossed 3 times.
Probability of getting NO heads on all three toss is 1/2*1/2*1/2 = 1/8
Therefore probability of getting at least 1 head is = 1 - 1/8 = 7/8